OpenStudy (anonymous):

can someone help me and find helpful tips with 9th grade coordinate algebra?

OpenStudy (anonymous):

Possibly, what's your question?

OpenStudy (anonymous):

i just need some help on that subject, like for examole the topic im on is writing and evaluating expressions

OpenStudy (anonymous):

or components of an algebraic expression

OpenStudy (anonymous):

Hmmm. Do you have any examples of questions? I'm not sure what you mean by coordinate algebra, but if I saw a problem, I'm sure I could help. There are lots of components of algebraic expressions, depending on the expression.

OpenStudy (anonymous):

example: susan is four years older than half nathans age. if nathan's age is represented by the variable n, what expression represents susans age?

OpenStudy (anonymous):

Sure, so in that one n=nathan's age. if susan is 4 years older, you'd have\[n+4=susan's~age\]You have a constant which is 4 and the variable, which is n. 4 is the same no matter what n is. So, n can be anything (technically. in this question it should be positive). Susan is always going to be 4 years older so the 4 won't change.

OpenStudy (anonymous):

You could write this\[f(n)=n+4\]no matter what you put in for n, you come out with Susan's age. Hopefully, that isn't getting too far ahead of things.

OpenStudy (anonymous):

oh ok, so all i have to do now is solve that equation and i get the answer!?

OpenStudy (anonymous):

Yep. You'd have a graph that looks like this

OpenStudy (anonymous):

Yep. All you do is enter an age for n and you get susan's age. So in that graph, you can see if Nathan is 0, Susan is 4. If Nathan is 3, Susan is 7.

OpenStudy (anonymous):

oh! ok i get it now, thanks for the help :)

OpenStudy (anonymous):

yep, no problem.

OpenStudy (phi):

just to tweak the answer: **susan is four years older than half nathans age.*** as you read this sentence, try to "toss out" irrelevant words. Look for numbers and "unknowns" translate "is" to = translate more than as + susan is four years older than half nathans age susan = 4 + n/2 notice that "half nathan's age" means nathan's age divided by 2. i.e. n/2 shorten susan to just a 1 letter variable that stands for susan's age s = 4 + n/2

OpenStudy (anonymous):

Oh goodness. Yep, I didn't read the half in the problem. Sorry @aryleen.tomlinson1234. Same idea, but phi has the correct answer to your problem.